OpenStudy (keke):

simplify; the square root of 216x^(2)

OpenStudy (anonymous):

VERY EASY 216x^2 = 36 * 6 * x^2 so root of 216x^2 = 6x * root of 6

OpenStudy (anonymous):

\[6\sqrt{6} x\]

OpenStudy (anonymous):

yes give me a medal